Can you factor 16c^6 - 2f^6
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

Can you factor 16c^6 - 2f^6
---------
Yes, I can.
----
Hint: Both terms have even coefficients.

Answer by Edwin McCravy(20060)

16c6 - 2f6
Factor out 2:

2(8c6 - f6)


Write the expression in parentheses as the difference of two
cubes.  The trickiest part is writing the first term 8c6 as a cube:

8c6 = 23c6 = 23(c2)3 = [2(c2)]3 = [2c2]3 = (2c2)3

f6) is easily written as a cube this way: (f2)3.

Change parentheses to brackets so you can have parentheses inside. 

 2[23c2)3 - (f2)3]

Rule: A³-B³ = (A-B)(A²+AB+B²)


2[(2c²) - (f²)][(2c²)² + (2c²)(f²) + (f²)²]

Sinplify

2(2c2-f2)(4c4+2c2f2+f4)


Edwin